LHC 2012 - latest news

Week 15: Machine coordinators: Mike Lamont, Bernhard Holzer

Plan for the week: safe intensity ramp up

 

Date

Summary

Link to report from morning meeting

Monday 16 April

        Power onverter problem, access needed

        06:30 prepare injection for VdM scans

8:30 meeting

Sunday 15 April

  • 01:15h dump, precycle & refill
  • 04:47h kicker failure during injection: MKID.B2, too short kicker pulse   vacuum valves closed, cryo lost, a number of quenches
  • experts called for check
  • 09:15 precycle to prepare new injection.
  • 13:15 stable beams
  • 23:00 scheduled dump

9:00 meeting

Saturday 14 April

  • 01:30h injection 1092x1092
  • 03:10h stable beams
  • 13:40 dump, precycle & re-fill
  • 16:14h luminosity

9:00 meeting

Friday 13 April

  • 07:05 Prepare for access.
  • 14:00h access finished, RF fault (Klystron filament)
  • 16:00h restart, injection 1092 bunches
  • 17:50 stable beams
  • 19:00h beam dump: ROD.A12.B1 (power conv: internal fault)

8:30 meeting

Thursday 12 April

  • 02:00 beam dump: elec. perturbation
  • 04:00 injection problems: pfn not ok
  • 07:30 injection for 840 bunches
  • 12:30 steering finished
  • First attempt with 840 bunches lost at end of squeeze to inadvertant turn on of ion pump in dump line.
  • 2 reasonable fills with 840 bunches. First lost to electrical network perturbation, second to main quad circuit S78 power converter.

8:30 meeting

Wednesday 11 April

  • 01:45 #2500 lost to controls problem (fibre optics repeater) which impacted cryogenics
  • Pre-cycle, re-injecting around 6:15
  • 08:00 Stable beams fill 2501 - still with 624 bunches - Alice solenoid and dipole off.
  • 10:44 beam dump due to (fake) BLM alarm in CMS
  • 13:45 injection for physics: 624 bunches
  • 18:00 stable beams

8:30 meeting

Tuesday 10 April

  • Overnight Monday to Tuesday: lost fill in ramp (orbit feedback problem); lost fill going into collisions (loss in 7, too high bunch intensity(?);
  • 04:00 Stable beams fill 2497 624 bunches - lost to trip of RCBXH1.R5 at 06:48
  • Access required for QPS for above trip, collimation and TCDQ teams in
  • Lift out in point 6 - repair required.
  • Unfortunately secondary collimator stuck in point 3 - will block out until next technical stop - loss maps required to check that things are still clean without it.
  • Beam back in 15:00 after pre-cycle
  • 2 loss maps at 4 TeV
  • Stable beam at around 20:00 - fill 2500

8:30 meeting - summary of last week

Monday 9 April

  • 4:50 Ramping down ALICE solenoid and dipole (+compensators)
  • 7:32 beams dumped during squeeze by BLM at TCLA.D6L7 RS 1.3 sec (losses on B2).
  • 10:30 Fill 2493 - Stable beams 624 bunches (ALICE dipole and solenoid off)
  • 16:54 Filll lost to trip of power converter RCBXV2.R1 - access required
  • 21:15 - looks like a flash-over of a beam 2 injection kicker - extended softstart required